Let the probability that the electrocardiograph will require repair during the warranty



period is 0.2. Determine the probability that during the warranty period out of six cardiographs:



1) no more than one will require repair; 2) at least one will require repair

1
Expert's answer
2022-01-17T16:14:21-0500

"n=6 \\

p = 0.2 \\

q = 1 -0.2 = 0.8 \\

P(X=x)=C(n,x)p^x q^{n-x}"

1) no more than one will require repair

"P(X≤1) = P(X=0) + P(X=1) \\

= \frac{6!}{0!(10-0)!} \times 0.2^0 \times 0.8^{6-0} + \frac{6!}{1!(6-1)!} \times 0.2^1 \times 0.8^{6-1} \\

= 0.262144 + 0.393216 \\

= 0.65536"

2) at least one will require repair

"P(X≥1) = 1 -P(X<1) \\

= 1 -P(X=0) \\

= 1 - [\frac{6!}{0!(10-0)!} \times 0.2^0 \times 0.8^{6-0}] \\

= 1 -0.262144 \\

= 0.737856"